# Two dimensional de-Sitter and deformed CFTs

### v2

@@ 1 Introduction
However, in this modest approach, the questions regarding the observer dependent choice of UV time is still obscure, since our description of worldsheet CFT already knows the ‘time’ of what observer will see in a static patch. We know in dS, different patches describe different notion of time and hence finding a global vacuum is inherently a difficult problem [^7] See [10] and the discussion in [21] from algebra of observables point of view. However, our present understanding might be more aligned with [31].. We end our note with some comments on it in section 6 along with possible future direction.
@@ 6 Discussion
In this note, our main goal is to find possible UV description of effective 2D theory of dilatons arising from reduction of higher dimensional Einstein gravity describing pure and asymptotically dS. While a stable dS2 vacuum is difficult to find in any time-dependent patches of dS2, we have managed to describe a sensible description of (extended and not pure) static patch effective theory in terms of worldsheet CFT with deformed Hamiltonian. Modular quantization of the worldsheet CFT further introduces a stretched horizon with an emergent conformal boundary condition, which acts like an internal observer in this description. In the CFT description, we have not taken $L\rightarrow iL$, since the role reversal of space-time direction in the proper time frame of the deformed Hamiltonian already manifests a CFT on 2d dS metric upto an overall change in Lorentzian signature. One may alternatively think of this set-up as de-Sitter mimicker, rather than an inherently de-Sitter. However, we would like to think of these deformed Hamiltonians of CFT generates 2d CFT on de-Sitter patches which, to the best of our knowledge, is not discussed previously in the literature. It would be definitely interesting to find a set-up where negatively divergent energy density in Milne vacuum could be compensated by gluing it with a very high positive energy density state like large AdS2 blackhole. This could probably lead to some version of cosmology in open universe.
@@
We end our discussion with some speculative remarks on the emergence of different worldsheet times.
@@ 6.1 Comment on emergence of observer time
Here $H_{F}$ is the time independent Floquet or effective Hamiltonian and generates a new stroboscopic time $t$. In this driven CFT set up, one can changes different phases dynamically by changing the frequency and amplitude of the drive parameter. By choosing appropriately $H_{1,2}$ one could end up with $H_{F}$ as in (6.1) [59], [60],[61]. Hence different stable configuration of $\alpha,\beta,\gamma$ should correspond to taking different phases. A detailed study of stability of those dynamical phases could be useful to understand those degenerate limits of charged SdS or other near extremal limits in AdS or flat black holes. For instance, from the phase diagram of [61], one can see the heating phases  [^37] This includes $\alpha^{2}<4\gamma^{2}$ as well as $\alpha=\gamma=0$ are separated from distinct non-heating phases by a phase boundary [^38] This includes $\alpha^{2}=4\beta^{2}$$\gamma=0$..
Nevertheless, the main point we want to make here is that the time $t$ inherent to the Floquet Hamiltonian $H_{F}$, is a stroboscopic time $t=n(T_{1}+T_{2})$. In this Floquet sense, the stroboscopic time is emergent. To be more precise, the Floquet observer will measure stroboscopic time in their own clock which is completely different than the clocks for the observers associated to $H_{1,2}$. This is analogous to find a static patch in an accelerating or time dependent (dS) universe. Consequently, to account for the cost of residing in an accelerating universe, static observers are required to observe features like surrounding cosmological (and black hole) horizons. We know that Floquet Hamiltonians in heating phases also exhibit features such as fixed points. We feel, the gravitational observers measure the stroboscopic Floquet time in the worldsheet description, which is inherently emergent. The principle behind the fine tuning of choosing the Hamiltonians $H_{1,2}$, drive parameters and frequencies still remains elusive to us from a gravitational perspective. We hope to return with more evidences on this in near future.
@@
For this metric $\bar{R}_{2}=-\frac{2}{L_{2}^{2}}$. Hence $I_{0}$ reduces to free massless scalar field action in the AdS2 Poincare background. $I_{2}$ is the topological term $I_{top}$ and it provides the extremal entropy of the 4D black hole. $I_{3}$ is the volume term which can be dropped by using suitable counter term. The boundary term $I_{1}$ should also be regulated by using a counter term. Hence the final form reduces to

### v3

@@ 1 Introduction
However, in this modest approach, the questions regarding the observer dependent choice of UV time is still obscure, since our description of worldsheet CFT already knows the ‘time’ of what observer will see in a static patch. We know in dS, different patches describe different notion of time and hence finding a global vacuum is inherently a difficult problem [^7] See [10] and the discussion in [21] from algebra of observables point of view. However, our present understanding might be more aligned with [31], as well as [10].. We end our note with some comments on it in section 6 along with possible future direction.
@@ 6 Discussion
In this note, our main goal is to find possible UV description of effective 2D theory of dilatons arising from reduction of higher dimensional Einstein gravity describing pure and black holes in dS. While a stable dS2 vacuum is difficult to find in any time-dependent patches of dS2, we have managed to describe a sensible description of (Nariai) static patch effective theory in terms of worldsheet CFT with deformed Hamiltonian. Modular quantization of the worldsheet CFT further introduces a stretched horizon with an emergent conformal boundary condition, which acts like an internal observer in this description. In the CFT description, we have not taken $L\rightarrow iL$, since the role reversal of space-time direction in the proper time frame of the deformed Hamiltonian already manifests a CFT on 2d dS metric upto an overall change in Lorentzian signature. To be more precise, the effective metric defined by the deformed CFT on strip has the following form upon analytic continuation in Lorentzian time:
$ds^{2}_{deformed}=4L^{2}_{AdS}\gamma^{2}\frac{ds^{2}-d\theta^{2}}{\sinh^{2}{2\gamma\theta}}=-\left(4L^{2}_{AdS}\gamma^{2}\frac{-ds^{2}+d\theta^{2}}{\sinh^{2}{2\gamma\theta}}\right)$ (6.1)
If we do not ignore the overall negative sign of the metric, one may interpret it as AdS2 Rindler patch with $\theta$ plays the role of time. On the other hand, if we could ignore the overall sign, we should think the same as Milne patch in de-Sitter [^32] Note that, if we wick-rotate the UHP coordinates $z\rightarrow iz$ and $\bar{z}\rightarrow i\bar{z}$, we end up with right half plane(RHP). The metric in the RHP is EAdS2 Poincare with an overall negative sign, which alternatively could be obtained by $L\rightarrow iL$ without wick-rotating the coordinates. Hence we may think of going UHP to RHP is same as changing $L\rightarrow iL$.. In the latter case, one may alternatively think of this set-up as de-Sitter mimicker, rather than an inherently de-Sitter due to negative cosmological constant [^33] This also justifies a negative stress tensor.. However, we would like to think of these deformed Hamiltonians of CFT generates 2d CFT on de-Sitter patches, which, to the best of our knowledge, is not discussed previously in the literature. It would be definitely interesting to find a set-up where negatively divergent energy density in our proposed Milne vacuum could be compensated by gluing it with a very high positive energy density state like large AdS2 blackhole. This could probably lead to some version of cosmology in open universe.
@@
We end our discussion with some remarks on the emergence of different worldsheet times.
@@ 6.1 Comment on emergence of observer time
Note that, taking $\gamma=0$ and $\alpha^{2}=4\beta^{2}$ would describe CFT in AdS2 Poincare(cold black hole), $\beta=0$ and $\alpha^{2}<4\gamma^{2}$ would describe CFT in $dS_{2}$ static patch and $\alpha=\gamma=0$ will describe modular Hamiltonian of a CFT on an interval of a ring[65]. It is well known that under (global) conformal transformation the modular flow generated by Rindler Hamiltonian is mapped to a spherical domain of a Lorentzian CFT[74]. For general $\alpha,\beta,\gamma\neq 0$, it would describe some complex metric which is not understood at this stage. Now we consider a slightly modified class of deformed Hamiltonian of a CFT on a ring of the form:
$H_{F}=i\alpha L_{0}+\beta(L_{1}+L_{-1})+i\gamma(l_{1}-L_{-1})+cc$ (6.3)
[…]
@@
Here $H_{F}$ is the time independent Floquet or effective Hamiltonian and generates a new stroboscopic time $t=n(T_{1}+T_{2})$. In this driven CFT set up, one can change different phases dynamically by changing the frequency and amplitude of the drive parameter. Hence the question in our context is whether we could naturally obtain a Floquet Hamiltonian of the form (6.3) by choosing certain $H_{1},H_{2}$.
Let us consider $H_{1}=a(L_{1}+L_{-1})$ and $H_{2}=ib(L_{1}-L_{-1})$ with real $(a,b)$. These are Hermitian operators as in radial quantization. By using Baker-Campbell-Hausdorff(BCH) formula we can in principle compute $e^{-iHT_{1}}e^{-iHT_{2}}$ as a nested commutators of $H_{1},H_{2}$ in Euclidean time. One could easily check that, for a single period this would take the form of (6.3) or $i\alpha L_{0}+\beta(L_{1}+L_{-1})+i\gamma(L_{1}-L_{-1})+cc$. Here $\alpha,\beta,\gamma$ depends of $a,b,c,T_{1},T_{2}$. Hence this apparent non-unitarity of the Floquet Hamiltonian arises from a Unitary periodically driven CFT set-up. However, in the Floquet time $t$, one can treat this as a Hermitian Hamiltonian with a different notion of Hermiticity [^40] The notion of Hermiticity explicitly depends on the choice of quantization time as in [67],[30]. The notion of adjoint operation in radial quantization may not coincide with a different quantization if the Hamiltonians are not unitarily related to each other..
Nevertheless, the main point we want to make here is that the time $t$ inherent to the Floquet Hamiltonian $H_{F}$, is a stroboscopic time. In this Floquet sense, the stroboscopic time is emergent. To be more precise, the Floquet observer will measure stroboscopic time in their own clock which is completely different than the clocks for the observers associated to $H_{1,2}$. This is analogous to find a static patch in an accelerating or time dependent (dS) universe. Consequently, to account for the cost of residing in an accelerating universe, static observers are required to observe features like surrounding cosmological (and black hole) horizons. We know that Floquet Hamiltonians in heating phases also exhibit features such as fixed points. We feel, the gravitational observers measure the stroboscopic Floquet time in the worldsheet description, which is inherently emergent. The principle behind the fine tuning of choosing the Hamiltonians $H_{1,2}$, drive parameters and frequencies still remains elusive to us from a gravitational perspective.
@@
For this metric $\bar{R}_{2}=-\frac{2}{L_{2}^{2}}$. Hence $I_{0}$ reduces to free massless scalar field action in the AdS2 Poincare background. $I_{2}$ is the topological term $I_{top}$ and it provides the extremal entropy of the 4D black hole. $I_{3}$ is the world-sheet volume term which may be dropped by using suitable counter term [^41] Note that, this is not required to drop this term. However, here our motivation is just to connect with JT gravity for which this may not be important to us.. The boundary term $I_{1}$ should also be regulated by using a similar counter term. Hence the final form reduces to